44^(2x-9 )+23=55
Could you solve please.

Answer by Alan3354(69443)

44^(2x-9 )+23=55
---
Subtract 23
44^(2x-9) = 32
---
(2x-9)*log(44) = log(32)
Can you do the rest?
------------
2x-9 = log(32)/log(44)
2x = 9 + log(32)/log(44)
x = 9/2 + log(32)/(2log(44))